A radioisotope has an initial activity of 5 millicurie. 48hours later, the observed activity is 4 millicurie. Determine: a) The half life of the radioisotope

b) The initial number of nuclides in the radioisotope.

Nb,Show workings



1
Expert's answer
2021-04-19T17:10:44-0400

"\displaystyle A(t) = A_0 2^{-\frac{t}{T_{1/2}}}"

a) "A_0= 5, A=4, t=48 h"

"\displaystyle \frac{A}{A_0} = 2^{-\frac{t}{T}}"

"\displaystyle \frac{t}{T}= log_2{\frac{A_0}{A}}=log_2 \frac{5}{4} = 0.3219"

"\displaystyle T = \frac{48 h}{0.3219}=149.11 \, h"


b) We will need to convert everything to SI units, curie -> Bq, h -> s.

"\displaystyle A_0= N_0 \frac{ln2}{T}"

"\displaystyle N_0 = \frac{A_0 T}{ln 2}=\frac{5 \cdot 10^{-3}\cdot 3.7⋅10^{10} \cdot 149.11 \cdot 3600}{0.6931} =14. 328 \cdot 10^{13}= 1.4328 \cdot 10^{14}"